What mass of water at 78.0 °C would we need to mix with 35.0 g of water at 14.5 °C to obtain a final water temperature of 19.5 °C ?
1
Expert's answer
2018-04-30T13:44:09-0400
The heat balance equation
m_1 c(t_1-t_3 )=m_2 c(t_3-t_2 )
Thus
m_1=(m_2 c(t_3-t_2 ))/(t_1-t_3 )=35.0(19.5-14.5)/(78.0-19.5)=3.00 g